# WARP-LCA: Efficient Convolutional Sparse Coding with Locally Competitive Algorithm

### v1

@@ 6 Conclusion
This research introduces WARP-LCA, an enhanced version of the Locally Competitive Algorithm (LCA), and demonstrates its efficacy in terms of accelerating convergence, improving robustness, and enabling the algorithm to achieve deeper minima and hence better solutions. As compared to the original LCA, we consistently observed improvements of image reconstruction metrics across various noise levels, and improvements of noise robustness across several deep neural network backbone architectures.
Our investigations further revealed the robust generalizability of WARP-LCA, particularly with the Oxford Pets dataset, which significantly differed from the training dataset (CIFAR) in image size, type, and image statistics.
@@
Looking ahead, expanding the capabilities of WARP-LCA to directly process noisy inputs could dramatically enhance its applicability for denoising, robustness, and other sparse coding applications.
Appendix A Parameters for Training LCA on CIFAR
Table 2: LCA hyperparameters on CIFAR-10.
Hyperparameter Value
Output neurons 100
Input neurons 3
Kernel size 9
Stride 2
Lambda 2.55
Tau 100
Eta 0.01
Lca iters 800
Pad same
Nonneg True
Transfer func soft_threshold
Figure 6: An exemplary dictionary of kernels obtained from training on the the CIFAR dataset
Appendix B WARP-CNN Architecture and Training hyper-parameters
Table 3: WARP-CNN Architecture
Layer Type Parameters Output Shape Common Layers Conv1 Conv2d 4, 512, kernel_size=5, stride=1, padding=2 (batch_size, 512, H, W) BN1 BatchNorm2d 512 (batch_size, 512, H, W) Conv2 Conv2d 512, 512, kernel_size=5, stride=1, padding=2 (batch_size, 512, H, W) BN2 BatchNorm2d 512 (batch_size, 512, H, W) Conv3 Conv2d 512, 512, kernel_size=3, stride=1, padding=1 (batch_size, 512, H, W) BN3 BatchNorm2d 512 (batch_size, 512, H, W) Conv4 Conv2d 512, 512, kernel_size=3, stride=1, padding=1 (batch_size, 512, H, W) BN4 BatchNorm2d 512 (batch_size, 512, H, W) Dropout Dropout p=0.3 (batch_size, 512, H, W) Conv5 Conv2d 512, 512, kernel_size=3, stride=1, padding=1 (batch_size, 512, H, W) BN5 BatchNorm2d 512 (batch_size, 512, H, W) Conv6 Conv2d 512, 512, kernel_size=3, stride=1, padding=1 (batch_size, 512, H, W) BN6 BatchNorm2d 512 (batch_size, 512, H, W) AdjustChannels Conv2d 512, 256, kernel_size=1 (batch_size, 256, H, W) Downward Branch Conv_d1 Conv2d 512, 256, kernel_size=3, stride=1, padding=1 (batch_size, 256, H, W) BN_d1 BatchNorm2d 256 (batch_size, 256, H, W) Conv_d2 Conv2d 256, 256, kernel_size=3, stride=1, padding=1 (batch_size, 256, H, W) BN_d2 BatchNorm2d 256 (batch_size, 256, H, W) Conv_d3 Conv2d 256, 100, kernel_size=3, stride=2, padding=1 (batch_size, 100, H/2, W/2) BN_d3 BatchNorm2d 100 (batch_size, 100, H/2, W/2) Upward Branch Conv_u1 Conv2d 512, 256, kernel_size=3, stride=1, padding=1 (batch_size, 256, H, W) BN_u1 BatchNorm2d 256 (batch_size, 256, H, W) Conv_u2 Conv2d 256, 256, kernel_size=3, stride=1, padding=1 (batch_size, 256, H, W) BN_u2 BatchNorm2d 256 (batch_size, 256, H, W) Conv_u3 Conv2d 256, 100, kernel_size=3, stride=2, padding=1 (batch_size, 100, H/2, W/2) BN_u3 BatchNorm2d 100 (batch_size, 100, H/2, W/2)
Table 4: Forward Pass of WARP CNN with Skip Connections
Step Operation Input Output Shape 1 Conv1 + ReLU x (batch_size, 512, H, W) 2 BN1 Output of Step 1 (batch_size, 512, H, W) 3 Conv2 + ReLU Output of Step 2 (batch_size, 512, H, W) 4 BN2 Output of Step 3 (batch_size, 512, H, W) 5 Conv3 + ReLU Output of Step 4 (batch_size, 512, H, W) 6 BN3 Output of Step 5 (batch_size, 512, H, W) 7 Conv4 + ReLU Output of Step 6 (batch_size, 512, H, W) 8 BN4 Output of Step 7 (batch_size, 512, H, W) 9 Dropout Output of Step 8 (batch_size, 512, H, W) 10 Conv5 + ReLU Output of Step 9 (batch_size, 512, H, W) 11 BN5 Output of Step 10 (batch_size, 512, H, W) 12 Dropout Output of Step 11 (batch_size, 512, H, W) 13 Conv6 + ReLU Output of Step 12 (batch_size, 512, H, W) 14 BN6 Output of Step 13 (batch_size, 512, H, W) 15 Dropout Output of Step 14 (batch_size, 512, H, W) 16 AdjustChannels Output of Step 15 (batch_size, 256, H, W) Downward Branch 17 Conv_d1 + ReLU Output of Step 15 (batch_size, 256, H, W) 18 BN_d1 Output of Step 17 (batch_size, 256, H, W) 19 Conv_d2 + ReLU Output of Step 18 + Output of Step 16 (batch_size, 256, H, W) 20 BN_d2 Output of Step 19 (batch_size, 256, H, W) 21 Conv_d3 + ReLU Output of Step 20 (batch_size, FEATURES, H/2, W/2) 22 BN_d3 Output of Step 21 (batch_size, FEATURES, H/2, W/2) Upward Branch 23 Conv_u1 + ReLU Output of Step 15 (batch_size, 256, H, W) 24 BN_u1 Output of Step 23 (batch_size, 256, H, W) 25 Conv_u2 + ReLU Output of Step 24 + Output of Step 16 (batch_size, 256, H, W) 26 BN_u2 Output of Step 25 (batch_size, 256, H, W) 27 Conv_u3 + ReLU Output of Step 26 (batch_size, FEATURES, H/2, W/2) 28 BN_u3 Output of Step 27 (batch_size, FEATURES, H/2, W/2) 29 Subtraction Output of Step 22 - Output of Step 28 (batch_size, FEATURES, H/2, W/2) 30 Sigmoid Output of Step 29 (batch_size, FEATURES, H/2, W/2)
Table 5: Training and testing details for the CIFAR-10 dataset.
Detail Description
Dataset CIFAR-10 standard training and test splits
Data Transformation ToTensor() Normalize(mean=[0.4914, 0.4822, 0.4465], std=[0.247, 0.243, 0.261])
Network Initialization net.apply(lambda m: sparse_init(m, sparsity=0.9, std=0.01))
Sparse Initialization Function def sparse_init(m, sparsity=0.9, std=0.01): if isinstance(m, (nn.Linear, nn.Conv2d)):     with torch.no_grad():      sparse_weights = torch.randn(m.weight.size()) * std      mask = torch.rand(m.weight.size()) > sparsity      sparse_weights = sparse_weights * mask.float()      m.weight = nn.Parameter(sparse_weights)      if m.bias is not None:       m.bias.data.zero_()
Optimizer Adam Learning Rate = 0.0001 optimizer = torch.optim.Adam(net.parameters(), lr=0.0001)
Loss Function custom_weighted_mse_loss_with_laplace_scale with non-zero weight scaling and Laplace-like scaling parameters. Non-zero weight: 1000.0, Laplace scale: 3.0.
Scaling of Target Activations Target activations are scaled between 0 and 1 during training, and predicted targets are descaled during inference.
Validation Data 1% of the training data was used for validation
Appendix C Training procedure for WARP-LCA
Algorithm 1 Training Dictionary, Encoding Data and Training CNN for WARP-LCA
1:Input: Dataset $D$, Number of epochs $E$, Initial $\lambda$, Increase factor $f$, Number of ISTA steps $n$
2:Output: Trained Dictionary $K$, Encoded Dataset $D^{\prime}$, Trained CNN
3:Initialize dictionary $K$
4:for $epoch=1$ to $E$ do
5:    for each sample in $D$ do
6:         Perform $n$ steps of ISTA with L1 sparsity (soft thresholding)
7:    end for
8:    $\lambda\leftarrow\lambda\times f$triangleright Increase sparsity factor
9:end for
10:Encode dataset:
11:for each sample in $D$ do
12:    Run LCA with hard thresholding
13:    Save activations to new dataset $D^{\prime}$
14:end for
15:Train CNN:
16:Initialize CNN
17:for each (input, label) in ($D$, $D^{\prime}$do
18:    Train CNN on input to predict corresponding label
19:end for
Appendix D Training on Tiny ImageNet
Figure 7: Encoding Performance Comparison between WARP-LCA and LCA The figure displays the performance metrics for WARP-LCA and LCA at $\lambda=0.15$. The columns represent different metrics: Mean Squared Error (MSE), L0 Norm, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM). For each plot, WARP-LCA and LCA performances are shown for comparison.
Table 6: Evaluation after 500 iterations on Tiny ImageNet dataset with $\lambda=0.15$
Metric Net Metrics Mean LCA Metrics Mean
MSE 0.0353 0.0588
$L_{0}$ Norm 6946.22 8169.27
$L_{1}$ Norm 3411.67 2710.40
PSNR 29.46 27.38
SSIM 0.9377 0.9092
Appendix E Comparing WARP-LCA and LCA at higher sparsity levels
Figure 8: Encoding Performance Comparison between WARP-LCA and LCA across Different Sparsity Levels. The figure displays the performance metrics for WARP-LCA and LCA over two sparsity levels, $\lambda=0.2$ and $\lambda=0.85$. Each row corresponds to a different sparsity level. The first row represents $\lambda=0.2$, while the second row represents $\lambda=0.85$. The columns represent different metrics: Mean Squared Error (MSE), L0 Norm, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM). For each plot, WARP-LCA and LCA performances are shown for comparison.
Appendix F More Examples for Accumulated Aggregation Maps
Figure 9: Larger Example of as shown in 5
Figure 10: Larger Example of as shown in 5. Both WARP-LCA and LCA were executed for 1000 iterations to fully converge. Visibly, the LCA image has a worse reconstruction than the WARP-LCA image. It seems to struggle with very dark or bright areas of the original image as can be seen seen in the difference plot in the second row (subtraction of original image and respective LCA method).
Figure 11: More Examples for Accumulated Aggregation Maps
Appendix G Datasets
Table 7: Details of the datasets used in the study.
Dataset License Source Authors Link
CIFAR-10 MIT License Canadian Institute For Advanced Research Alex Krizhevsky, Geoffrey Hinton https://www.cs.toronto.edu/~kriz/cifar.html
Oxford Pets Creative Commons Attribution 4.0 International License University of Oxford O. M. Parkhi, A. Vedaldi, A. Zisserman https://www.robots.ox.ac.uk/~vgg/data/pets/
Tiny ImageNet Public Domain Stanford University Li-Jia Li, Fei-Fei Li https://huggingface.co/datasets/zh-plus/tiny-imagenet
Appendix H Equivalence of LCA with Hard Thresholding to IHTA and Distinction from Matching Pursuit
H.1 Equivalence to IHTA with Hard Thresholding
We follow the example of the soft thresholding equivalence, given by Balavoine et al. [2015], to establish the equivalence between the Iterative Hard Thresholding Algorithm (IHTA) and LCA with hard thresholding, consider the IHTA update rule:
$a^{(k+1)}=H_{k}\left(a^{(k)}+\eta\Phi^{T}(y-\Phi a^{(k)})\right)$ (7)
where $H_{k}(\cdot)$ is the hard thresholding operator that retains the top $k$ largest (in magnitude) elements and sets the rest to zero.
Rewriting the LCA ODE with a discretization step size $\Delta l=\tau$:
$\tau\frac{u[k+1]-u[k]}{\tau}=-u[k]+\Phi^{T}(y-\Phi a[k])-(\Phi^{T}\Phi-I)a[k]$ (8)
Assuming $u[k]=a[k]$ for the initial state, we have:
$u[k+1]=a[k]+\Phi^{T}(y-\Phi a[k])-(\Phi^{T}\Phi a[k]-a[k])$ (9)
Simplifying this:
$u[k+1]=a[k]+\Phi^{T}(y-\Phi a[k])$ (10)
Then apply the hard thresholding operator:
$a[k+1]=H_{k}(u[k+1])$ (11)
Comparing this with the IHTA update:
$a^{(k+1)}=H_{k}\left(a^{(k)}+\eta\Phi^{T}(y-\Phi a^{(k)})\right)$ (12)
This shows that when assuming $u[k]=a[k]$ for the initial state, the update for $u[k+1]$ and applying the hard thresholding operator $H_{k}$ leads to the same update rule as IHTA. Therefore, LCA with hard thresholding is equivalent to IHTA when the initial state assumption $u[k]=a[k]$ is used.
H.2 Distinction from Matching Pursuit with Hard Thresholding
Matching Pursuit (MP) employs a greedy, sequential strategy that iteratively selects the dictionary element with the highest correlation to the residual. The steps of MP are:
  1. 1.
    Initialize the residual $r^{0}=y$ and set $a^{0}=0$.
  2. 2.
    For each iteration $k$:
    • Find the column $\Phi_{j}$ that best matches the residual: $j=\arg\max_{i}|\langle r^{k},\Phi_{i}\rangle|$.
    • Update the coefficients: $a_{j}^{k+1}=a_{j}^{k}+\langle r^{k},\Phi_{j}\rangle$.
    • Update the residual: $r^{k+1}=r^{k}-\langle r^{k},\Phi_{j}\rangle\Phi_{j}$.
LCA with hard thresholding ($\ell_{0}$ regularization) uses a parallel approach where all neurons compete simultaneously. The competition term $(I-\Phi^{T}\Phi)a^{(k)}$ in LCA introduces dynamics absent in MP. Thus, LCA with $\ell_{0}$ regularization is not equivalent to Matching Pursuit.

### v2

@@ 6 Conclusion
This research introduces WARP-LCA, an enhanced version of the Locally Competitive Algorithm (LCA), and demonstrates its efficacy in terms of accelerating convergence, improving robustness to noise and the initial choice of parameters, and enabling the algorithm to achieve deeper minima and hence better solutions. As compared to the original LCA, we consistently observed improvements of image reconstruction metrics across various noise levels, and improvements of noise robustness across several deep neural network backbone architectures.
Our investigations further revealed the robust generalizability of WARP-LCA, particularly with the Oxford Pets dataset, which significantly differed from the training dataset (CIFAR) in image size, type, and image statistics. Moreover, WARP-LCA proved highly effective in initializing other thresholding algorithms, such as CELO and L1/2, leading to substantially improved outcomes. By providing a strong initial estimate that aligns well with the underlying data structure, WARP-LCA facilitates faster convergence and enhances the precision of these subsequent methods, yielding significantly better results in practice. Additionally, we found that a good initialization—such as that provided by WARP-CNN—permits the use of a larger step size during optimization, thereby enabling even faster convergence while maintaining stability and accuracy.
@@
Looking ahead, expanding the capabilities of WARP-LCA to directly process noisy inputs could dramatically enhance its applicability for denoising, robustness, and other sparse coding applications. Furthermore, alternative approaches—such as specialized application-tailored models or deep sparse coding architectures—could be employed to complement or extend the current framework, offering additional pathways for improved performance and flexibility in diverse scenarios.